# An effective Bombieri–Vinogradov error term for sifting problems

### v1

@@ 2.3  Bounds relating to $\pi(x;d,a)$
Essentially, we show (Proposition (2.14)) that one can obtain an effective Bombieri–Vinogradov type theorem for (2.27) whenever $k\mid k_{1}$ with $k\neq 1$.
@@
where $\chi(n)$ is a Dirichlet character. Here, one has the relation (see e.g. [24, p. 377])
@@
Let $d\leq Q_{1}$ and $\beta_{d}$ denote the Siegel zero modulo $d$ if it exists. Then for any $y_{0}\leq y\leq\mathcal{X}$ and $a\in\mathbb{Z}$ with $(a,d)=1$,
$\left|\sum_{\begin{subarray}{c}\chi\ (\mathrm{mod}\ d)\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\psi(y,\chi)\right|\ll_{B}\begin{cases}\frac{y}{(\log y)^{B}}+y^{\beta_{d}},&\text{if}\ \beta_{d}\ \text{exists},\\ \frac{y}{(\log y)^{B}},&\text{otherwise,}\end{cases}$ (2.29)
where the implied constant is effective and independent of $a$.
@@
From (2.28), we have
$\left|\sum_{\begin{subarray}{c}\chi\ (\mathrm{mod}\ d)\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\psi(y,\chi)\right|\leq\varphi(d)\left|\psi(y;d,a)-\frac{\psi(y,\chi_{0})}{\varphi(d)}\right|,$
@@
$\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}\mu^{2}(d)\gamma^{\omega(d)}\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,kd)=1}\left|\pi(y;kd,a)-\frac{\pi(y;k,a)}{\varphi(d)}\right|\ll_{\gamma,B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B_{\gamma}}}.$ (2.31)
@@
The proof is similar to [4, Lemma 4.10] but we give the full details for completeness. First we prove the corresponding result for the Chebyshev $\psi$ function and $\gamma=1$. That is,
$\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,kd)=1}\left|\psi(y;kd,a)-\frac{\psi(y;k,a)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}.$ (2.32)
Let $y\leq\operatorname{\mathcal{X}}$ and $d,k,a\in\mathbb{N}$ such that $(d,\overline{\mathcal{P}})=1$, $(d,k)=1$ and $(a,kd)=1$. Now, since $(k,d)=1$ any character modulo $kd$ can be represented in a unique way as the product of two characters modulo $k$ and modulo $d$. Thus,
@@
$\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}\frac{\mu^{2}(d)}{\varphi(kd)}\left|\sum_{\begin{subarray}{c}\chi_{1}\>(\text{mod }k)\\ \chi_{2}\neq\chi_{0,d}\>(\text{mod }d)\end{subarray}}\overline{\chi_{1}}(a)\overline{\chi_{2}}(a)\psi(y;\chi_{1}\chi_{2})\right|\ll\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-2}}.$ (2.35)
However, this can be proven in an identical fashion to Proposition 2.8, using Lemma 2.13 as opposed to Lemma 2.5. Here we also do not need the restriction $k_{1}\leq\log\operatorname{\mathcal{X}}$ since $\chi_{2}\neq\chi_{0,d}$ and $k\mid k_{1}$, so that the exceptional character never appears in the inner sum of (2.35).
Now, to convert this to a result in terms of the prime counting function $\pi$ and arbitrary $\gamma\geq 0$, one applies partial summation as in the proof of Proposition 2.10, followed by the Cauchy–Schwarz inequality as in the proof of Proposition 2.12, giving (2.31). ∎
@@ 3  Proof of the effective upper bound
$V_{1}(z)\leq\left(1+O_{A}\left(\frac{1}{\log\log\operatorname{\mathcal{X}}}\right)\right)V(z).$
@@
Applying the conditions (1.15) then gives
@@
Thus, if we can show that $q_{1}\gg\log\log\operatorname{\mathcal{X}}$ then we are done. In this direction, first note that since $k_{1}$ is square-free, so too is $q_{1}$. Thus, since $k_{1}\geq\log\operatorname{\mathcal{X}}$ we have $q_{1}\geq\overline{p}$ where $\overline{p}$ is the largest prime such that
$\log\operatorname{\mathcal{X}}\geq\prod_{p\leq\overline{p}}p,$
or by taking logarithms,
$\log\log\operatorname{\mathcal{X}}\geq\theta(\overline{p}),$
where $\theta(x)=\sum_{p\leq x}\log p$ is the Chebyshev theta prime-counting function. So in particular,
$\theta(q_{1})\geq\log\log\operatorname{\mathcal{X}}.$
The desired bound $q_{1}\gg\log\log\operatorname{\mathcal{X}}$ then follows by the prime number theorem, or even the Brun–Titchmarsh theorem (Lemma 2.6). ∎
@@
Case 1: $k_{1}\leq\log\operatorname{\mathcal{X}}$. In this case, the result follows directly from (2.21) of Proposition 2.12.
@@ 4.1  Preliminary lemmas
To begin with, we use an inclusion-exclusion argument to relate $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ to sifting functions in terms of $\operatorname{\mathcal{A}}_{m_{j}}$ and $\operatorname{\mathcal{P}}_{j}$. This is the key identity that will allow us to circumvent the Siegel zero in the sieving error term. This lemma is essentially the same as [4, Lemma 6.3], but we include the short proof here for completeness.
@@
with $m_{j}$ and $\operatorname{\mathcal{P}}_{j}$ as defined in (2.5).
@@
First we note that
$V_{1}(z)-V(z)=\frac{V(z)}{1-g(q_{1})}-V(z)=\frac{g(q_{1})V(z)}{1-g(q_{1})}.$
Thus,
$g(m_{j})V_{j+1}(z)=\prod_{i=1}^{j}{g(q_{i})}\frac{V(z)}{\prod_{i=1}^{j+1}(1-g(q_{i}))}=\left(V_{1}(z)-V(z)\right)\frac{\prod_{i=2}^{j}g(q_{i})}{\prod_{i=2}^{j+1}(1-g(q_{i}))}$ (4.5)
Applying the bounds $g(q_{i})\leq A/(q_{i}-1)$ and $q_{i}\geq A+2$ to (4.5) then gives
$g(m_{j})V_{j+1}(z)$ $\leq\left(V_{1}(z)-V(z)\right)\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1)\prod_{i=2}^{j+1}(1-A/(q_{i}-1))}$
$=(V_{1}(z)-V(z))\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1)}\left(1+\frac{A}{q_{j+1}-1-A}\right)$
$\leq(V_{1}(z)-V(z))\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1)}\left(1+A\right)$
$\leq 2A(V_{1}(z)-V(z))\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1)}.$ (4.6)
Since each $q_{i}$ is distinct and $q_{i}\geq A+2$, (4.6) can then further be bounded as
$g(m_{j})V_{j+1}(z)\leq 2A(V_{1}(z)-V(z))\frac{A^{j-1}}{(j-1)!}.$ (4.7)
By similar reasoning,
$g(m_{j})V_{\ell}(z)\leq(V_{1}(z)-V(z))\frac{A^{\ell-1}}{(\ell-1)!}<2A(V_{1}(z)-V(z))\frac{A^{\ell-1}}{(\ell-1)!}.$ (4.8)
Combining (4.7) and (4.8) then gives
$\sum_{j=1}^{\ell-1}g(m_{j})V_{j+1}(z)+g(m_{\ell})V_{\ell}(z)\leq 2A(V_{1}(z)-V(z))\sum_{j=1}^{\ell}\frac{A^{j-1}}{(j-1)!}\leq 2Ae^{A}(V_{1}(z)-V(z)),$
as required. ∎
@@
$r(m_{j})\ll_{B,\gamma,\varepsilon}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1-\varepsilon}}.$
@@
$\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,m_{j})=1}\left|\pi(y;m_{j},a)-\frac{\pi(y)}{\varphi(m_{j})}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-1}}.$
Here, since $j<\ell$ we note that $m_{j}$ does not divide the exceptional modulus up to $Q_{1}(\operatorname{\mathcal{X}})=(\log\operatorname{\mathcal{X}})^{B}$, so $\beta(m_{j})$ (appearing in (2.8)) is bounded by Lemma 2.3. To finish, we use the bound $m_{j}<k_{1}\leq(\log\operatorname{\mathcal{X}})^{B}$ as well as (4.9) of Lemma 4.4 to obtain
@@
The case $y\leq\sqrt{\operatorname{\mathcal{X}}}$ follows from the trivial bounds $\pi(y;k_{1},a)\leq y$ and $\pi(y)\leq y$, so we assume $y\geq\sqrt{\operatorname{\mathcal{X}}}$. By the Brun–Titchmarsh theorem (Lemma 2.6) and the fact that $k_{1}\leq(\log\operatorname{\mathcal{X}})^{B}$},
@@
$E_{j}:=\begin{cases}\sum_{d<D_{j},\>(d,\overline{\operatorname{\mathcal{P}}}_{j+1})=1}\gamma^{\omega(d)}\mu^{2}(d)|r_{m_{j}}(d)|,&\text{if $j=0,\ldots,\ell-1$},\\ \sum_{d<D_{\ell},\>(d,\overline{\operatorname{\mathcal{P}}}_{\ell})=1}\gamma^{\omega(d)}\mu^{2}(d)|r_{m_{\ell}}(d)|,&\text{if $j=\ell$}.\end{cases}$
Applying the upper bound sieve (1.12) gives,
@@
noting that $|\operatorname{\mathcal{A}}|=X$. Here, by the condition ($\mathbf{\Omega}(\kappa,L)$), we have
@@
$\qquad\geq XV(z)+V(z)\cdot O_{L,B,\gamma,\varepsilon}\left((\log\log\operatorname{\mathcal{X}})^{\kappa}\left(\frac{\operatorname{\mathcal{X}}(\log\log\operatorname{\mathcal{X}})}{(\log\operatorname{\mathcal{X}})^{B-1-\varepsilon}\log\log\log\operatorname{\mathcal{X}}}+\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{2-\varepsilon}}\right)\right)$
@@
where in the last step we have suitably rescaled $\varepsilon$, and used that $B>3$ and $\operatorname{\mathcal{X}}\asymp X\log X$. This completes our analysis of the first line of (4.16).
@@
$\sum_{j=0}^{\ell}E_{j}\leq\omega(k_{1})\cdot O_{\gamma,B}\left(\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-\gamma-1}}\right)\ll_{\gamma,B}\frac{\operatorname{\mathcal{X}}\log\log\log\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-\gamma-1}\log\log\operatorname{\mathcal{X}}}.$ (4.28)
@@
$\Pi_{2}(x)$ $=\pi(z)+S(\operatorname{\mathcal{A}}_{1},\operatorname{\mathcal{P}}_{1},z),$ (5.1)
@@
Moreover, via an application of Merten’s third theorem,
@@
Thus, in the context of Theorem 1.6, we have a sifting problem with $A=1$ and $\kappa=1$. We now use the Rosser–Iwaniec linear sieve upper bound [13, Theorem 1]
@@
and for $s>3$, $F(s)$ is defined by a differential-delay equation and satisfies the conditions of Theorem 1.6, see [9, §8.2]. Here, we note that the $O_{L}$ term in (5.4) is effective and a fully explicit variant is given in [4, Theorem 2.1]. Thus, letting
$D=\frac{x}{(\log x)^{4}},$
we can apply Theorem 1.6 with $\gamma=1$, $B=4$, $s=1$ and $V(z)$ given in (5.3) to obtain the effective upper bound
@@
$<4C_{2}\frac{\pi(x)}{\log x}\left(1+O\left(\frac{1}{\log\log X}\right)\right)+\pi(\sqrt{x})+O\left(\frac{\pi(x)}{(\log x)^{2}}\right).$
The desired bound (1.4) then follows upon applying an effective and suitably strong version of the prime number theorem $\pi(x)\sim x/\log x$ (e.g. [31, Theorem 1]). ∎
@@
It is also common to use Selberg’s upper bound sieve [9, Theorem 3.2] to obtain an upper bound for $\Pi_{2}(x)$ or $G(n)$. Selberg’s sieve would give a simpler $O_{L}$ term in (5.4) yet a slightly more complicated remainder term. Ultimately, using either sieve (Rosser-Iwaniec or Selberg) yields the same upper bound asymptotically, but the size of $X(\varepsilon)$ and $N(\varepsilon)$ in Theorem 1.2 would differ.
@@
We note here that any $a\in\operatorname{\mathcal{A}}$ has no divisors in $\overline{\operatorname{\mathcal{P}}}$. To see this, consider some $a=n-q^{2}\in\operatorname{\mathcal{A}}$. Then $(a,n)=1$ since $(q,n)=1$. In addition, $2\nmid a$ since $n\equiv 0\pmod{2}$} and $q^{2}\equiv 1\pmod{2}$. Similarly, $3\nmid a$. All this is to say is that $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ gives a lower bound for the number of elements of $\operatorname{\mathcal{A}}$ with no prime factor less than $z$. In particular, for the purposes of the proof, we wish to show that
@@
For $n$ to give a solution to (5.5) we require that $n$ is a quadratic residue modulo $p$. That is, there exists some $a_{0}\pmod{p}$} such that $a_{0}^{2}\equiv n$ (mod $p$). If this is the case, then there are two congruence classes of solutions to (5.5), namely $a_{0}$ and $a_{0}$. In particular,
@@
noting that in (5.6) we used that there are $O(\log n)$ unique prime divisors of $n$. By the Chinese remainder theorem, the asymptotic (5.8) extends to
@@
for any square-free $d$ with $(d,\overline{\mathcal{P}})=1$. Thus, for our sifting problem it is suitable to define

### v2

@@ 2.3  Bounds relating to $\pi(x;d,a)$
Essentially, we show (Proposition 2.13) that one can obtain an effective Bombieri–Vinogradov type theorem for (2.27) whenever $k\mid k_{1}$ with $k\neq 1$.
@@
where $\chi$ is a Dirichlet character. Here, one has the relation (see e.g. [25, p. 377])
@@
Let $d\leq Q_{1}$ and $\beta_{d}$ denote the Siegel zero modulo $d$ if it exists. Then,
$\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,d)=1}\left|\sum_{\begin{subarray}{c}\chi\ (\mathrm{mod}\ d)\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\psi(y,\chi)\right|\ll_{B}\begin{cases}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}}+\operatorname{\mathcal{X}}^{\beta_{d}},&\text{if}\ \beta_{d}\ \text{exists},\\ \frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B}},&\text{otherwise,}\end{cases}$ (2.29)
where the implied constant is effective.
@@
Let $y\leq\operatorname{\mathcal{X}}$ and $(a,d)=1$. From (2.28), we have
$\left|\sum_{\begin{subarray}{c}\chi\ (\mathrm{mod}\ d)\\ \chi\neq\chi_{0}\end{subarray}}\overline{\chi}(a)\psi(y,\chi)\right|=\varphi(d)\left|\psi(y;d,a)-\frac{\psi(y,\chi_{0})}{\varphi(d)}\right|,$
@@
${\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}}^{\!\!\!\!\!\sharp}\ \mu^{2}(d)\gamma^{\omega(d)}\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,kd)=1}\left|\pi(y;kd,a)-\frac{\pi(y;k,a)}{\varphi(d)}\right|\ll_{B,\gamma}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B_{\gamma}}},$ (2.31)
where $\sharp$ means that the sum is restricted to $d$ with $k_{1}\nmid kd$ if $k\neq k_{1}$.
@@
The proof is similar to [3, Lemma 4.10] but we give the full details for completeness. We also remark that in [3, Lemma 4.10] and its proof there are a couple of small errors, particularly the lack of the $\sharp$ restriction on the moduli $d$. Thus, we offer a corrected statement and proof here.
First we prove the corresponding result for the Chebyshev $\psi$ function and $\gamma=1$. That is,
$\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}\mu^{2}(d)\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,kd)=1}\left|\psi(y;kd,a)-\frac{\psi(y;k,a)}{\varphi(d)}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-3}}.$ (2.32)
Let $y\leq\operatorname{\mathcal{X}}$ and $d,k,a\in\mathbb{N}$ such that $(d,\overline{\mathcal{P}})=1$, $(d,k)=1$ and $(a,kd)=1$. Since $(k,d)=1$ any character modulo $kd$ can be represented in a unique way as the product of two characters modulo $k$ and modulo $d$. Therefore,
@@
$\sum_{\begin{subarray}{c}d\leq D/k\\ (d,\overline{\operatorname{\mathcal{P}}})=1\\ (d,k)=1\end{subarray}}\frac{\mu^{2}(d)}{\varphi(kd)}\left|\sum_{\begin{subarray}{c}\chi_{1}\>(\text{mod }k)\\ \chi_{2}\neq\chi_{0,d}\>(\text{mod }d)\end{subarray}}\overline{\chi_{1}}(a)\overline{\chi_{2}}(a)\psi(y;\chi_{1}\chi_{2})\right|\ll\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B-3}}.$ (2.35)
However, this can be proven in an identical fashion to Proposition 2.7, using Lemma 2.12 as opposed to Lemma 2.5. Here we also do not need the restriction $k_{1}\leq\log\operatorname{\mathcal{X}}$ since the conditions $\chi_{2}\neq\chi_{0,d}$, $k\mid k_{1}$ and the $\sharp$ restriction on the moduli $d$ ensure that the exceptional character never appears in the inner sum of (2.35).
To convert this to a result in terms of the prime counting function $\pi$ and arbitrary $\gamma\geq 0$}, one uses partial summation as in the proof of Proposition 2.9, followed by the Cauchy–Schwarz inequality as in the proof of Proposition 2.11, giving (2.31). ∎
@@ 3  Proof of the effective upper bound
$V_{1}(z)\leq\left(1+O_{A}\left(\frac{1}{\log\log\operatorname{\mathcal{X}}}\right)\right)V(z).$ (3.1)
@@
Applying the condition $g(q_{1})\leq A/(q_{1}-1)$ (see (1.16)) then gives
@@
Thus, if we can show that $q_{1}\gg\log\log\operatorname{\mathcal{X}}$ then we are done. In this direction, first recall that $k_{1}$ is square-free by its definition (2.4). Thus,
$\prod_{p\leq q_{1}}p\geq k_{1}>\log\operatorname{\mathcal{X}}.$
Taking logarithms,
$\theta(q_{1})>\log\log\operatorname{\mathcal{X}},$
where $\theta(x)=\sum_{p\leq x}\log p$ is the Chebyshev theta prime-counting function. The desired bound $q_{1}\gg\log\log\operatorname{\mathcal{X}}$ then follows by the prime number theorem. ∎
@@
Case 1: $k_{1}\leq\log\operatorname{\mathcal{X}}$. Here, the result follows from (2.21) of Proposition 2.11.
@@ 4.1  Preliminary lemmas
To begin with, we use an inclusion-exclusion argument to relate $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ to sifting functions in terms of $\operatorname{\mathcal{A}}_{m_{j}}$ and $\operatorname{\mathcal{P}}_{j}$. This is the key identity that will allow us to circumvent the Siegel zero in the sieving error term. This lemma is essentially the same as [3, Lemma 6.3], but we include the short proof here for completeness.
@@
with $m_{j}$ and $\operatorname{\mathcal{P}}_{j}$ as defined in (2.5) and (2.6) respectively.
@@
We begin by extending our notation for $q_{i}$ (defined in Section 2.1) by setting
$q_{i}=0,\qquad i>\ell$
and use the convention that $g(0)=0$. Hence, with this notation
$V_{\ell+1}(z)=\frac{V_{\ell}(z)}{1-g(q_{\ell+1})}=V_{\ell}(z).$
By further noting that (see Lemma 3.1)
$V_{1}(z)-V(z)\ll_{A}V(z),$
we see that it is sufficient to prove
$\sum_{j=1}^{\ell}g(m_{j})V_{j+1}(z)\ll_{A}V_{1}(z)-V(z)$ (4.6)
in place of (4.5). Next, we note that
$V_{1}(z)-V(z)=\frac{V(z)}{1-g(q_{1})}-V(z)=\frac{g(q_{1})V(z)}{1-g(q_{1})}$
and thus,
$g(m_{j})V_{j+1}(z)$ $=\prod_{i=1}^{j}{g(q_{i})}\frac{V(z)}{\prod_{i=1}^{j+1}(1-g(q_{i}))}$
$=\frac{V_{1}(z)-V(z)}{1-g(q_{i+1})}\prod_{i=2}^{j}\frac{g(q_{i})}{1-g(q_{i})}.$ (4.7)
In order to bound (4.7), we split into two cases $1\leq j<j^{\prime}$ and $j\geq j^{\prime}$, where $j^{\prime}$ is defined to be the largest integer with
$q_{i}\geq A+2\quad\text{for all}\ i\in\{1,\ldots,j^{\prime}\},$ (4.8)
meaning also that there are only $O_{A}(1)$ values of $j$ with $j^{\prime}\leq j\leq\ell$.
Case 1: $1\leq j<j^{\prime}$.
In this case, we substitute the bound $g(q_{i})\leq A/(q_{i}-1)$ from (1.16) and the bound (4.8) into (4.7) to yield
$g(m_{j})V_{j+1}(z)$ $\leq\frac{V_{1}(z)-V(z)}{1-A/(q_{j+1}-1)}\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1-A)}$
$\ll_{A}(V_{1}(z)-V(z))\frac{A^{j-1}}{\prod_{i=2}^{j}(q_{i}-1-A)}.$ (4.9)
Since each $q_{i}$ in the product in (4.9) is distinct and satisfies $q_{i}\geq A+2$ by (4.8), (4.9) can be bounded further as
$g(m_{j})V_{j+1}(z)\ll_{A}(V_{1}(z)-V(z))\frac{A^{j-1}}{(j-1)!}.$ (4.10)
Case 2: $j\geq j^{\prime}$.
In this case, we first note that since there are only $O_{A}(1)$ values of $j$ with $j^{\prime}<j\leq\ell$,
$\prod_{j>j^{\prime}}\frac{1}{1-g(q_{j})}=O_{A}(1).$
Therefore, from (4.7) and the fact that $g(q_{i})<1$, we have for $j\geq j^{\prime}$,
$g(m_{j})V_{j+1}(z)$ $\ll_{A}(V_{1}(z)-V(z))\prod_{i=2}^{j^{\prime}}\frac{g(q_{i})}{1-g(q_{i})}$
$\ll(V_{1}(z)-V(z))\frac{A^{j^{\prime}-1}}{(j^{\prime}-1)!}$
with the second bound following from the argument in Case 1.
Combining Cases 1 and 2,
$\sum_{j=1}^{\ell}g(m_{j})V_{j+1}(z)\ll_{A}(V_{1}(z)-V(z))\left(\sum_{1\leq j<j^{\prime}}\frac{A^{j-1}}{(j-1)!}+\sum_{j^{\prime}\leq j\leq\ell}\frac{A^{j^{\prime}-1}}{(j^{\prime}-1)!}\right).$
Here,
$\sum_{1\leq j<j^{\prime}}\frac{A^{j-1}}{(j-1)!}\leq e^{A}$
and since (j1)!(j-1)! grows faster than $A^{j-1}$,
$\sum_{j^{\prime}\leq j\leq\ell}\frac{A^{j^{\prime}-1}}{(j^{\prime}-1)!}\ll_{A}\sum_{j^{\prime}\leq j\leq\ell}1=O_{A}(1).$
Thus, (4.6) and thereby (4.5) follows. ∎
@@
$r(m_{j})\ll_{B,\gamma,\varepsilon}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B+1-\varepsilon}}.$
@@
$\sup_{y\leq\operatorname{\mathcal{X}}}\max_{(a,m_{j})=1}\left|\pi(y;m_{j},a)-\frac{\pi(y)}{\varphi(m_{j})}\right|\ll_{B}\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B+1}}.$
Here, since $j<\ell$ we note that $m_{j}$ is not divisible the exceptional modulus up to $Q_{1}(\operatorname{\mathcal{X}})=(\log\operatorname{\mathcal{X}})^{B}$, so $\operatorname{\mathcal{X}}^{\beta_{m_{j}}}$ (appearing in (2.10)) is suitably bounded by Lemma 2.2. To finish, we use the bound $m_{j}<k_{1}\leq(\log\operatorname{\mathcal{X}})^{B}$ as well as (4.11) of Lemma 4.4 to obtain
@@
The case $y\leq\sqrt{\operatorname{\mathcal{X}}}$ follows from the trivial bounds $\pi(y;k_{1},a)\leq y$ and $\pi(y)\leq y$}, so we assume $y\geq\sqrt{\operatorname{\mathcal{X}}}$. Then, by the Brun–Titchmarsh theorem,
@@
$E_{j}:=\begin{cases}\sum_{d<D_{j},\>(d,\overline{\operatorname{\mathcal{P}}}_{j+1})=1}\gamma^{\omega(d)}\mu^{2}(d)|r_{m_{j}}(d)|,&\text{if $j=0,\ldots,\ell-1$},\\ \sum_{d<D_{\ell},\>(d,\overline{\operatorname{\mathcal{P}}}_{\ell})=1}\gamma^{\omega(d)}\mu^{2}(d)|r_{m_{\ell}}(d)|,&\text{if $j=\ell$},\end{cases}$
with $\operatorname{\mathcal{P}}_{j}$ as defined in (2.6). Applying the upper bound sieve (1.13) gives,
@@
noting that $|\operatorname{\mathcal{A}}|=X$. Here, by the condition (1.12), we have
@@
$\qquad\geq XV(z)+V(z)\cdot O_{L,B,\gamma,\varepsilon}\left((\log\log\operatorname{\mathcal{X}})^{\kappa}\left(\frac{\operatorname{\mathcal{X}}(\log\log\operatorname{\mathcal{X}})}{(\log\operatorname{\mathcal{X}})^{B+1-\varepsilon}\log\log\log\operatorname{\mathcal{X}}}+\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{2-\varepsilon}}\right)\right)$
@@
where in the last step we have suitably rescaled $\varepsilon$, and used that $B>1$ and $\operatorname{\mathcal{X}}\asymp X\log X$. This completes our analysis of the first line of (4.18).
@@
$\sum_{j=0}^{\ell}E_{j}\leq\omega(k_{1})\cdot O_{\gamma,B}\left(\frac{\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B_{\gamma}}}\right)$ $\ll_{B,\gamma}\frac{\operatorname{\mathcal{X}}\log\log\log\operatorname{\mathcal{X}}}{(\log\operatorname{\mathcal{X}})^{B_{\gamma}}\log\log\operatorname{\mathcal{X}}}$
$\ll\frac{X\log\log\log X}{(\log X)^{B_{\gamma}-1}\log\log X}$ (4.30)
@@
$\Pi_{2}(x)$ $\leq\pi(z)+S(\operatorname{\mathcal{A}}_{1},\operatorname{\mathcal{P}}_{1},z),$ (5.1)
@@
Moreover, via an application of Mertens’ theorem,
@@
Thus, in the context of Theorem 1.6, we have a sifting problem with $A=1$ and $\kappa=1$. We now use the Rosser–Iwaniec linear sieve upper bound [15, Theorem 1]
@@
and for $s>3$, $F(s)$ is defined by a differential-delay equation and satisfies the conditions of Theorem 1.6, see [11, §8.2]. Here, we note that the $O_{L}$ term in (5.4) is effective and a fully explicit variant is given in [3, Theorem 2.1]. Thus, letting
$D=\frac{\sqrt{x}}{(\log x)^{5}},$
we can apply Theorem 1.6 with $A=1$, $\gamma=1$, $B=5$, $s=1$ and $V(z)$ given in (5.3) to obtain the effective upper bound
@@
$<\pi(\sqrt{x})+4C_{2}\frac{\pi(x)}{\log x}\left(1+O\left(\frac{1}{\log\log X}\right)\right)+O\left(\frac{\pi(x)}{(\log x)^{2}}\right).$
The desired bound (1.4) then follows upon applying an effective version of the prime number theorem $\pi(x)\sim x/\log x$ (e.g. [33, Theorem 1]). ∎
@@
It is also common to use Selberg’s upper bound sieve [11, Theorem 3.2] to obtain an upper bound for $\Pi_{2}(x)$ or $G(n)$. Selberg’s sieve would give a simpler $O_{L}$ term in (5.4) yet a slightly more complicated remainder term. Ultimately, using either sieve (Rosser-Iwaniec or Selberg) yields the same upper bound asymptotically, but the size of $X(\varepsilon)$ and $N(\varepsilon)$ in Theorem 1.2 would differ.
@@
We first confirm that any $a\in\operatorname{\mathcal{A}}$ has no divisors in $\overline{\operatorname{\mathcal{P}}}$. To see this, consider some $a=n-q^{2}\in\operatorname{\mathcal{A}}$. Then $(a,n)=1$ since $(q,n)=1$. In addition, $2\nmid a$ since $n\equiv 0\pmod{2}$} and $q^{2}\equiv 1\pmod{2}$. Similarly $3\nmid a$. All this is to say is that $S(\operatorname{\mathcal{A}},\operatorname{\mathcal{P}},z)$ gives a lower bound for the number of elements of $\operatorname{\mathcal{A}}$ with no prime factor less than $z$. In particular, for the purposes of the proof, we wish to show
@@
For $n$ to give a solution to (5.5) we require that $n$ is a quadratic residue modulo $p$. That is, there exists $a_{0}\pmod{p}$} with $a_{0}^{2}\equiv n$ (mod $p$). If this is the case, then there are two congruence classes of solutions to (5.5), namely $a_{0}$ and $a_{0}$. In particular,
@@
where in (5.6) we used that there are $O(\log n)$ unique prime divisors of $n$. By the Chinese remainder theorem, the asymptotic (5.8) extends to
@@
for any square-free $d$ with $(d,\overline{\mathcal{P}})=1$. Thus, for our sifting problem we set